A wire 1.0m long and with cross sectional area 2.0×10^-7m^2 has a resistance of 0.1Ω calculate the electrical conductivity of the wire
1
Expert's answer
2020-05-29T10:09:23-0400

"L=1.0m\\S=2\times10^{-7}m^2\\R=0.1\Omega\\\rho-?\\solution:\\R=\rho\times \frac{L}{S} \Rightarrow\rho=\frac{R\times S}{L};\\\rho=\frac{0.1\times2\times10^{-7} }{1}=0.2\times 10^{-7}\Omega\times m.\\Answer:\;the \;electrical\; conductivity\\ of \;the\; wire\;equals\;0.2\times 10^{-7}\Omega\times m."